OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (anonymous):

Let a be first term and r be common ratio, then tn=ar^(n-1) -12=ar^(2-1)=ar ...(1) 768=ar^(5-1) =ar^4 ....(2) divide (2) by (1) \[\frac{ ar ^{4} }{ar }=\frac{ 768 }{-12 },r ^{3}=-64=\left( -4 \right)^{3},r=-4\]

OpenStudy (anonymous):

a(-4)=-12 a=-12/-4 =3 \[tn=3\left( -4 \right)^{n-1}\]

OpenStudy (anonymous):

It's okay, I already figured that out. Thank You :)